Given the following data, determine the molecule.
Homework #3 M
Mass Spec Homework Problem #3 m/z 14 15 16 intensity 11 52 -- 25 2 - W
Answer #1

The molecular ion peak (M+) corresponds to m/z = 88

There is a peak at m/z = 73

88-73 = 15 corresponds to CH3-fragment.

There is a peak at m/z = 56

73-56 = 17 corresponds to OH-fragment.

The base peak at m/z = 43 corresponds to the propyl fragment (CH3CH2CH2).

The peak at m/z = 29 corresponds to the CH3-CH2 fragment.

Therefore, the structure of the molecule can be written as follows.

CH3-CH2-CH2-CH(OH)-CH3